The height, y, of a ball tossed into the air can be represented by the equation y=-x^2+10x+3, where x is the elasped time. How many seconds does it take the ball to reach a height of 28 feet?

Answer by nerdybill(7384)

The height, y, of a ball tossed into the air can be represented by the equation y=-x^2+10x+3, where x is the elasped time. How many seconds does it take the ball to reach a height of 28 feet?
.
set y to 28 and solve for x:
y=-x^2+10x+3
28=-x^2+10x+3
0=-x^2+10x-25
0=x^2-10x+25
0=(x-5)(x-5)
x = 5 seconds